OpenStudy (anonymous):

The area of a rectangle is 44ft squared, and the length of the rectangle is 3 ft less than twice the width. find the dimensions of the rectangle.

OpenStudy (anonymous):

Area=L*w L=2w-3 AREA=(2W-3)*W 44=2w^2-3w when you factor it out you get -4 (which isnt a possible length) and 11/2 . so width is 11/2 and L = 2(11/2)-3 ..... which is length of 8 . to check:you do L*W=A. 8*11/2=44 CORRECT :)